$SPAD/src/input VECTOR.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- Input generated from VectorXmpPage
)clear all
u : VECTOR INT := new(5,12)
v : VECTOR INT := vector([1,2,3,4,5])
#(v)
v.2
v.3 := 99
v
5 * v
v * 7
w : VECTOR INT := vector([2,3,4,5,6])
v + w
v - w

———-

References^{[1] nothing}

4